OpenStudy (anonymous):

For an experiment, Mike plans to flip a coin and roll a fair, six-sided die at the same time. 1 2 3 4 5 6 Heads Tails Use the area model above to determine the probability that the coin will land on heads and the die will land on a number greater than 4.

OpenStudy (anonymous):

can someone help?

OpenStudy (anonymous):

\[\frac{1}{2} \times \frac{1}{6} + \frac{1}{2} \times \frac{1}{6} \]